1. Corporation reports that at an activity level of 9,300 units, its total variable cost is $669,600 and its total fixed cost is $96,650. What would be the total cost, both fixed and variable, at an activity level of 9,500 units? Assume that this level of activity is within the relevant range.

  • $766,250

  • $774,489

  • $780,650

  • $783,750

  • 2.The following costs were incurred in April:


  • Direct materials $48,000
    Direct labor $32,000
    Manufacturing overhead $26,000
    Selling expenses $23,000
    Administrative expenses $36,000


    Prime costs during the month totaled:

  • $80,000

  • $165,000

  • $58,000

  • $106,000

3.

alip Corporation, a merchandising company, reported the following results for October:

Sales $ 424,800
Cost of goods sold (all variable) $ 180,100
Total variable selling expense $ 18,200
Total fixed selling expense $ 15,300
Total variable administrative expense $ 8,700
Total fixed administrative expense $ 31,200


The contribution margin for October is:

  • $351,400

  • $217,800

  • $244,700

  • $171,300
Answer #1

1) Calculate total Cost

Variable cost (669600/9300*9500) 684000
Fixed cost 96650
Total 780650

So answer is c) $780650

2) Prime cost = Direct material+Direct labor = 48000+32000 = 80000

So answer is a) $80000

3) Contribution margin = 424800-180100-18200-8700 = 217800

So answer is b) $217800